OpenStudy (blowery):

Simplify this square root Square root 144x^9y^7 I think the next step is to take the square root of 144, so 12(x^9y^7)^1/2. What is next?

jimthompson5910 (jim_thompson5910):

divide 9 into 2 to get 9/2 = \(\Large {\color{red}{4}}\) remainder \(\Large {\color{blue}{1}}\) the \(\Large {\color{red}{4}}\) will be the exponent for the x outside the square root the \(\Large {\color{blue}{1}}\) will be the exponent for the x inside the square root so \[\Large \sqrt{x^9} = x^{{\color{red}{4}}}\sqrt{x^{{\color{blue}{1}}}} = x^4\sqrt{x}\]

jimthompson5910 (jim_thompson5910):

Another example. Let's say we had the square root of x^17 Divide 17 in half to get 17/2 = \(\Large {\color{red}{8}}\) remainder \(\Large {\color{blue}{1}}\) Which means \[\Large \sqrt{x^{17}} = x^{{\color{red}{8}}}\sqrt{x^{{\color{blue}{1}}}} = x^8\sqrt{x}\]

OpenStudy (blowery):

I wish I could write like that. So, would my answer be 12x^4 ( sq rt x) x^3 (sq rt x) 12 x^7 times x 12x^8 ?

jimthompson5910 (jim_thompson5910):

You forgot about y. There aren't two x terms

jimthompson5910 (jim_thompson5910):

But yes, \[\Large \sqrt{y^7} = y^3\sqrt{y}\]

OpenStudy (blowery):

Duh, let me try again

jimthompson5910 (jim_thompson5910):

btw you can use the equation editor (the blue "equation" button below the text box)

jimthompson5910 (jim_thompson5910):

or you can type it in by hand like so type in `\(\Large \sqrt{x}\)` to get \(\Large \sqrt{x}\)

OpenStudy (blowery):

12x^4(sq rt x)y^3(sq rt y)

jimthompson5910 (jim_thompson5910):

from there you can combine the stuff in the roots using this rule \[\Large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\]

OpenStudy (blowery):

I'm using an iPad and don't see the equation editor?

jimthompson5910 (jim_thompson5910):

hmm I'm not sure how the layout is for the iPad. I guess you'll have to type it in by hand if you want the equations to render like that

jimthompson5910 (jim_thompson5910):

Anyways, using that rule, you'll go from this \[\Large 12x^4\sqrt{x}*y^3\sqrt{y}\] to this \[\Large 12x^4y^3\sqrt{xy}\]

jimthompson5910 (jim_thompson5910):

So \[\Large \sqrt{144x^9y^7}\] fully simplifies to \[\Large 12x^4y^3\sqrt{xy}\]

OpenStudy (blowery):

I really appreciate your help!

jimthompson5910 (jim_thompson5910):

glad to be of help